# Comprehensive Review of Deep Unfolding Techniques for Next-Generation Wireless Communication Systems

### v3

@@ VII-A Discussions and Takeaways
In Table VII, we summarize the most popular unfolded network for the decoding of channel codes. Deep unfolding spans their applications in modern error-correcting codes, such as LDPC, turbo, and polar codes. The soft probabilistic estimates are progressively refined through learnable decoding layers in deep unfolding while enforcing parity constraints, enabling near-optimal hard decisions without the need for manual hyperparameter tuning. Hence, it merges model based optimization, for structured soft-output refinement and codeword validity with data-driven learning, for adaptive penalty tuning, outperforming traditional decoders in accuracy and efficiency.
@@ X-A Limitations of Deep Unfolding Techniques
Although deep unfolding has found applications in various design problems, it has limitations. Most of these limitations arise from the fact that deep unfolding is not an independent design methodology; it must be built upon existing conventional optimization frameworks. Consequently, it faces the following challenges in the development and real-world applications:
@@
(C1) Methodological challenges: Deep unfolding methods are inspired by the iterative procedures of conventional optimization algorithms, where each iteration is typically interpreted as a layer in an NN. In traditional wireless transmitter-receiver design and signal processing, optimization problems are often non-convex and highly coupled due to factors such as fractional SINR, constant-modulus constraints (e.g., for analog beamforming), or mixed-integer variables (e.g., in symbol detection or antenna selection). They are typically solved using a two-layer approach. In the first layer, AO or block coordinate ascent/descent is used to decouple variables, breaking the original problem into smaller, more manageable sub-problems. In the second layer, convex optimization techniques such as SCA or the ADMM are employed to solve each sub-problem. While this hierarchical structure enables the use of convex solvers (e.g., CVX), it often introduces auxiliary variables and results in nested loops that must converge at each layer, significantly increasing algorithmic complexity.
Deep unfolding models, on the other hand, aim to transform these iterative procedures into end-to-end trainable architectures. Instead of solving sub-problems sequentially, the unfolding framework typically approximates the entire optimization process in a more unified and parallelizable manner. This shift in design philosophy, from explicitly solving decoupled sub-problems to learning a unified mapping, poses non-trivial challenges, especially when attempting to translate complex multi-stage optimizers with auxiliary variables into a compact, learnable structure. Therefore, while deep unfolding builds upon conventional iterative methods, the methodological differences in structure and implementation represent a major challenge in developing effective unfolding models.
@@
(C2) Unfolding principle algorithms: While deep unfolding typically enhances conventional iterative algorithms, not all can be effectively unfolded. This limitation arises from the core principle of deep unfolding, which structures the iterations of an iterative algorithm as layers within a DNN. Complex operations like matrix decomposition or the use of convex optimization tools such as CVX or YALMIP (with solvers like MOSEK, SeDuMi, etc.) are difficult to implement in a DNN framework. In large-scale systems such as massive MIMO, even matrix multiplications result in high complexity for unfolding models.
Another challenge is parameter generalization, particularly with respect to the number of layers in the unfolding model. In iterative algorithms with nested loops (discussed in C1), both inner and outer loops run until convergence, but deep unfolding models require the number of layers to be predetermined before deployment. This is typically done through empirical tuning, which may not provide optimal values for all channel realizations. As a result, convergence is not guaranteed across different scenarios, potentially leading to performance degradation. Proper tuning is essential to achieve a favorable performance-complexity tradeoff in deep unfolding models.
(C3) Handling design constraints and objectives: One of the biggest challenges in developing deep unfolding models lies in managing the design constraints formulated in the problem. These constraints require the DNN layers’ outputs to meet specific requirements, such as total power, QoS, or other element-wise constraints. Simple constraints, like power or constant modulus, can often be enforced through normalization [181]. For instance, the $tanh$ activation function can be employed to impose constant-modulus constraints on analog beamforming coefficients or sensing waveforms, similar to black-box DL architectures. However, QoS constraints are significantly more complex. While they resemble activation functions, none of the existing activation functions can adequately manage such intricate requirements. One approach is to incorporate these constraints into the design objective. However, this technique often results in an nonequivalent problem with a more complicated objective function [191].
Not only do complex constraints pose challenges for developing deep unfolding models, but intricate objective functions also present difficulties in training the models. This is because their loss functions are generally formulated based on the objective function. For instance, unfolding architectures like DetNet [49], FSNet [75], and ScNet [229] for symbol detection are trained to minimize the Euclidean distance between the model’s output (symbol vector) and the true transmitted symbol vector. In particular, when labels are unavailable, as in beamforming problems, unfolding networks use the objective function as the loss function for minimization or its negative for maximization problems. Typical deep unfolding techniques [191, 155] require calculating the gradient of the objective function. However, this becomes challenging when the objective function is a complex, high-order, or fractional function of the design variables, especially when those variables are vectors or matrices.
(C4) Sophisticated mappings: While deep unfolding is powerful, certain inference tasks still require black-box DNN models, especially when there is no clear relationship between the input data, the output solution, and performance metrics. For instance, in multi-modal sensing-assisted beam tracking, tasks may involve complex mappings from environmental data, such as sensory information from sensors and cameras, to the optimal beam [230, 231]. In such cases, there is no explicit relationship between environmental images and the optimal beam, and standard communication metrics like throughput or SINR cannot be directly expressed as functions of these images. These problems cannot be formulated as optimization tasks solvable with principle algorithms, making it difficult to develop unfolding models that map environmental information to beam patterns or blockage states.
@@
(C5) Real-world applications: Deep unfolding architectures, while promising in mimicking iterative algorithms through trainable layers, often entail considerable depth and parameter complexity [191]. This imposes severe challenges for real-time inference under strict latency and throughput constraints, especially in systems with multiple and large dimensions like massive MIMO-OFDM. From a hardware perspective, deep unfolding designs are not naturally aligned with the requirements for area-efficient and parallelizable architectures needed in scalable OFDM deployments. For example, the non-uniform and sequential computations common in unfolding detectors can result in inefficient silicon utilization and high power consumption, limiting the ability to instantiate multiple parallel detection cores across subcarriers [232].
Furthermore, in real-world scenarios, the data distribution is highly dynamic because of user mobility, varying propagation environments, and hardware impairments. These variations often violate the assumptions underlying many unfolding-based transmitter and receiver designs, which are typically developed under idealized conditions. However, this assumption breaks down in practical deployments, where channels often exhibit strong line-of-sight components and spatial correlation. As a result, detectors relying on such assumptions can suffer significant performance degradation. To maintain robustness in these conditions, deep unfolding models must be redesigned or trained using data that captures the complexity and variability of realistic, deployment-specific channels.
The above five challenges (C1C5) also outline the scenarios wherein deep unfolding models cannot be built or does not offer much gain in performance-complexity tradeoff. These include when there is lack of suitable conventional iterative algorithms (C1, C2, C4), when the design problems are highly complicated with constraints and objective function (C3), or when the exist imperfection in real-world propagation environments and hardware (C5). It can be seen that deep unfolding is not a game changer that will completely replace conventional principle algorithms or black-box ML models. However, the applicability of deep unfolding can be enhanced by overcoming those challenges, as will be discussed next.
@@ X-B1 Coexistence of Deep Unfolding Models, Conventional Solvers, and Black-Box Models
Deep unfolding doesn’t need to fully replace conventional algorithms or black-box ML models. Instead, these approaches can coexist, leveraging their combined strengths. In hybrid designs, deep unfolding can (i) reduce the number of sub-problems, (ii) lower the complexity of solving each sub-problem, and (iii) enhance the learning process in black-box models. Benefits (i) and (ii) correspond to challenges C1 and C2, while (iii) relates to C4.
@@ X-B2 Deep Unfolding for Multi-Objective Learning
X-B3 Distributed Deep Unfolding Architectures
While deep unfolding models provide advantages in terms of complexity due to faster convergence compared to conventional iterative algorithms, they can still face high complexity issues due to challenge C2. This limits their applicability on platforms with constrained computational resources, such as edge devices or distributed access points (APs). Future wireless networks will advance beyond simple communications, evolving into large-scale, distributed systems such as cell-free massive MIMO that integrate advanced sensing, communications, and computing capabilities. However, most existing AI models, both data-driven and model-driven, were designed for centralized scenarios, where they are deployed at a central node with substantial computing power. These models are not well-suited for distributed scenarios, highlighting the need for appropriate distributed unfolding architectures.
One potential solution is to develop distributed deep unfolding architectures. In this approach, rather than executing all the unfolded layers at every distributed node or edge device, the layers can be distributed across multiple platforms for parallel execution, thereby reducing latency. For example, to design $L$ beamformers for $L$ APs for their downlink transmission in a distributed system, a CPU with high computing power and hardware resources can handle the majority of the layers, generating high-quality beamformers. These beamformers are then sent to the APs, serving as initial solutions and inputs for smaller sub-unfolding models at the distributed APs. This ensures that reliable beamformers can be generated at the APs with only a few layers, minimizing complexity and runtime to meet the computational constraints of the APs. Conversely, in the uplink, a distributed unfolding detection model can be deployed at the APs to generate initial detected symbols, which are then forwarded to the CPU for further processing and final decision-making by the large unfolding detector. By such an unfolding framework, highly reliable symbols are generated with low complexity at distributed APs.
@@ X-B4 Deep Unfolding to Enhance Black-box Solution
As its name suggests, a black-box ML model operates with a black-box nature, making it unexplainable and resistant to re-structuring beyond tuning parameters like the number of layers or nodes. The performance and efficiency of such models are heavily dependent on the quality of input data. Deep unfolding can contribute by performing pre-processing steps to improve the input data quality. One example is in multi-modal sensing-assisted communication tasks, where data from distributed cameras, LiDAR, and radar sensors are used to enhance communication performance, such as in beam tracking and blockage prediction. However, significant challenges arise in collecting, transferring, and processing large sensory data. Specifically, high-resolution camera images generate considerable traffic overhead and latency, impacting real-time communication processing. Deep unfolding models can be particularly useful in pre-processing such high resolution images. In image processing, deep unfolding has been successfully applied to tasks such as denoising, super-resolution, and image reconstruction, allowing for high-quality outcomes with fewer iterations than conventional algorithms. For example, unfolded algorithms based on ISTA for sparse coding have been shown to significantly accelerate image reconstruction while maintaining high-quality results [235, 236, 237]. This makes deep unfolding a valuable tool for handling large-scale sensory data in real-time communication systems.

### v4

@@ VII-A Discussions and Takeaways
Table VII compares various unfolded networks for decoding in error correcting codes based on computational complexity, trainable parameter count, and convergence speed (in number of layers). Deep unfolding spans their applications in modern error-correcting codes, such as LDPC, turbo, and polar codes. The soft probabilistic estimates are progressively refined through learnable decoding layers in deep unfolding while enforcing parity constraints, enabling near-optimal hard decisions without the need for manual hyperparameter tuning. Hence, it merges model based optimization, for structured soft-output refinement and codeword validity with data-driven learning, for adaptive penalty tuning, outperforming traditional decoders in accuracy and efficiency.
@@ X-A Limitations of Deep Unfolding Techniques
Although deep unfolding has found applications in various design problems, it also has important limitations. Most of these stem from the fact that deep unfolding is not an independent design methodology; it must be built upon existing conventional optimization frameworks. Consequently, it faces the following challenges in both development and real-world applications:
@@
(C1) Methodological challenges: Deep unfolding methods are inspired by the iterative procedures of conventional optimization algorithms, where each iteration is typically interpreted as a layer in a neural network. In traditional transmitter-receiver design and signal processing, optimization problems are often non-convex and highly coupled due to factors such as fractional SINR, constant-modulus constraints, or mixed-integer variables. These are typically solved by a two-layer approach: in the first layer, alternating optimization or block coordinate ascent/descent is used to decouple variables; in the second layer, convex optimization techniques such as SCA or ADMM are applied to the resulting sub-problems. While this makes the problems solvable using convex solvers (e.g., CVX), it introduces auxiliary variables and nested loops that must converge at each stage, increasing algorithmic complexity. Deep unfolding, by contrast, attempts to approximate such iterative processes in a unified, trainable architecture, making it challenging to translate multi-stage solvers into compact unfolding models.
@@
(C2) Unfolding principle algorithms: Deep unfolding typically enhances conventional iterative algorithms, but not all can be effectively unfolded. Complex operations such as matrix decompositions or convex solver calls (e.g., CVX, YALMIP with MOSEK or SeDuMi) are difficult to represent in differentiable NN layers. Furthermore, large-scale systems such as massive MIMO incur high complexity even for basic operations like matrix multiplications. Another issue is the fixed number of layers: iterative algorithms run until convergence, whereas unfolding models require the number of layers to be fixed a priori, often through empirical tuning. This can lead to suboptimal performance or lack of guaranteed convergence across different scenarios.
(C3) Handling design constraints and objectives: Another major challenge is enforcing problem-specific constraints and objectives in the unfolding framework. Simple constraints (e.g., power or constant modulus) can often be imposed by normalization or suitable activation functions, but more complex QoS constraints are difficult to embed. Similarly, intricate objective functions result in complicated loss functions, especially when gradients are hard to compute for high-order or fractional functions. This complicates training and may limit the range of problems amenable to unfolding.
(C4) Sophisticated mappings: While deep unfolding is powerful for problems with well-defined iterative solvers, many inference tasks lack such formulations. For example, in multi-modal sensing-assisted beam tracking, there is no explicit relationship between environmental sensory data (e.g., images) and the optimal beam. Since these problems cannot be expressed as tractable optimization tasks, unfolding methods are not suitable, and black-box learning models remain necessary.
@@
(C5) Real-world applications: Deep unfolding architectures, while promising in theory, often require considerable network depth and parameter complexity. This creates challenges for real-time inference under stringent latency and throughput constraints, especially in high-dimensional systems such as massive MIMO-OFDM. Hardware-wise, unfolding designs are not naturally aligned with area-efficient or parallelizable architectures, leading to inefficient silicon utilization and high power consumption. Moreover, practical deployments face dynamic channel conditions, mobility, and hardware impairments that often violate the assumptions made in unfolding-based designs. As a result, models trained under idealized conditions may suffer significant performance degradation in practice.
The above five challenges (C1C5) illustrate scenarios where deep unfolding is either not applicable or offers limited performance-complexity benefits. These include the absence of suitable iterative algorithms (C1, C2, C4), the presence of highly complicated constraints or objectives (C3), and practical imperfections in real-world propagation environments and hardware (C5). It is clear that deep unfolding is not a universal replacement for conventional algorithms or black-box machine learning models, but rather a complementary tool most effective when problem structures are well matched to unfolding principles.
At the same time, several strategies can mitigate the negative impacts of these limitations in practice. Robustness can be improved by training on realistic datasets that incorporate mobility, propagation variability, and hardware impairments. Hybrid designs, in which unfolding is applied only to specific modules (e.g., replacing matrix inversion or accelerating convergence) while other parts rely on conventional or data-driven methods, can achieve better tradeoffs between interpretability and performance. Finally, hardware-aware unfolding architectures tailored for parallelism and energy efficiency can help overcome deployment challenges in large-scale systems. Moreover, the applicability of deep unfolding can be enhanced by overcoming those challenges, as will be discussed next.
@@ X-B1 Coexistence of Deep Unfolding Models, Conventional Solvers, and Black-Box Models
Deep unfolding doesn’t need to fully replace conventional algorithms or black-box machine learning models. Instead, these approaches can coexist, leveraging their combined strengths. In hybrid designs, deep unfolding can (i) reduce the number of sub-problems, (ii) lower the complexity of solving each sub-problem, and (iii) enhance the learning process in black-box models. Benefits (i) and (ii) correspond to challenges C1 and C2, while (iii) relates to C4.
@@ X-B2 Deep Unfolding for Multi-Objective Learning
X-B3 Scalable Distributed Deep Unfolding Architectures
While deep unfolding models provide advantages in terms of complexity due to faster convergence compared to conventional iterative algorithms, they can still face scalability issues due to challenge C2, especially when deployed on platforms with constrained computational resources such as edge devices or distributed access points (APs). Future wireless networks will evolve into large-scale distributed systems, such as cell-free massive MIMO, that integrate advanced sensing, communications, and computing capabilities. However, most existing AI models, both data-driven and model-driven, are designed for centralized scenarios where they are executed at a powerful central node, making them poorly suited for distributed deployments. This highlights the need for scalable distributed unfolding architectures.
One potential solution is to develop distributed deep unfolding frameworks. Instead of executing all the unfolded layers at every AP or edge device, the layers can be partitioned across multiple platforms for parallel execution, thereby reducing latency and computational burden. For example, to design $L$ precoders for $L$ APs, a CPU with strong computational resources can process the majority of the layers and generate high-quality beamformers. These beamformers are then sent to the APs as initial solutions, which are refined locally by lightweight sub-unfolding models requiring only a few layers. This ensures reliable beamformers at the APs while maintaining scalability by reducing their computational load. Conversely, in the uplink, a distributed unfolding model can be deployed at the APs to generate initial solutions, which are then forwarded to the CPU for final refinement by a larger unfolding model.
Such a hierarchical CPU-AP cooperation framework directly addresses scalability challenges in distributed deployments by lowering the per-node computational cost, reducing fronthaul/backhaul signaling overhead, and enabling heterogeneous APs to operate with resource-aware lightweight sub-models. In practice, this can be further enhanced by tailoring the unfolding depth at each AP to its available hardware resources and by adopting split-learning style execution, where unfolding layers are adaptively divided between CPU and APs. In this way, reliable symbol detection and beamforming can be achieved with scalable complexity across distributed networks.
@@ X-B4 Deep Unfolding to Enhance Black-box Solution
As its name suggests, a black-box machine learning model operates with a black-box nature, making it unexplainable and resistant to re-structuring beyond tuning parameters like the number of layers or nodes. The performance and efficiency of such models are heavily dependent on the quality of input data. Deep unfolding can contribute by performing pre-processing steps to improve the input data quality. One example is in multi-modal sensing-assisted communication tasks, where data from distributed cameras, LiDAR, and radar sensors are used to enhance communication performance, such as in beam tracking and blockage prediction. However, significant challenges arise in collecting, transferring, and processing large sensory data. Specifically, high-resolution camera images generate considerable traffic overhead and latency, impacting real-time communication processing. Deep unfolding models can be particularly useful in pre-processing such high resolution images. In image processing, deep unfolding has been successfully applied to tasks such as denoising, super-resolution, and image reconstruction, allowing for high-quality outcomes with fewer iterations than conventional algorithms. For example, unfolded algorithms based on ISTA for sparse coding have been shown to significantly accelerate image reconstruction while maintaining high-quality results [46, 244, 230]. This makes deep unfolding a valuable tool for handling large-scale sensory data in real-time communication systems.
X-B5 Evaluation Protocols for Deep Unfolding
To facilitate the development of deep unfolding techniques, it is necessary to establish standardized evaluation protocols. A natural and broadly applicable criterion is the convergence profile, i.e., how quickly an unfolding model achieves satisfactory performance within a limited number of layers [152, 153, 132, 107, 156, 157, 149]. This metric reflects both computational complexity (since fewer layers correspond to fewer iterations and lower runtime) and model performance. Convergence profiles also allow direct comparison with conventional iterative algorithms, highlighting the performance complexity trade-offs of unfolding designs.
Another important aspect is the methodology for employing different learnable hyperparameters. Most unfolding techniques learn scalar parameters such as step sizes [152, 153, 132, 107, 156, 157, 149], while others consider richer structures such as trainable matrices [83]. Although the specific parameters vary, they share the common effect of shaping convergence speed and stability, which again makes the convergence profile a valid and unifying evaluation criterion. By contrast, training convergence criteria (e.g., number of epochs or early stopping) are less critical, as unfolding models typically converge quickly [152], and this detail is often omitted in the literature. Furthermore, reproducible open-source implementations of deep unfolding solutions with clear and consistent settings would not only enable fair comparisons but also facilitate the development of new unfolding models. In Table VIII, we present several available implementations for representative tasks.
TABLE VIII: Open-source implementations of deep unfolding techniques. The GitHub links are embedded directly in model names as hyperlinks.
Design Problem Representative Works
Theoretical convergence of LISTA LISTA [33]
Detection DetNet [163], FSNet [151], MMNet [101], LoRD-Net [102], CG-OAMP-Net [246]
Channel Estimation SBL-Net [54]
Precoding Unfolded WMMSE [156]
Hybrid Beamforming ManNet, Sub-ManNet [152], Unfolded PGA [107], Modular PGA [108]
ISAC Unfolded PGA-ISAC [153]
Decoding in Error Correcting Codes JCED-Net [180], NMS-Net [198]